# Position: Simulating Society Requires Simulating Thought

### v2

@@ 3.3 Metric Illusion: Current Benchmarks are Far from Enough
Despite the growing sophistication of generative agents, evaluation standards remain stuck in a prior era of NLP—optimizing for stylistic fluency, short-term coherence, and plausibility of individual outputs. Most benchmarks treat language as a proxy for thought, assuming that coherent expression implies coherent reasoning. As a result, current evaluations risk rewarding agents for each response’s local plausibility, ignoring whether they maintain internal consistency across turns or scenarios. Stance classification tasks, for example, check whether a model picks a side but remain agnostic about how or why that stance was formed [79, 80]. Dialogue benchmarks reward conversational smoothness, even when agents flip positions over time [81, 82, 83]. These limitations flatten reasoning to performance—encouraging shallow persona simulation without testing whether agents have structured beliefs that guide their answers.
More critically, today’s benchmarks rarely assess belief revision or causal adaptability. Current evaluations assess agents on static inputs but neglect their responses to counterfactual interventions or other hypothetical changes. When prompted with counterfactuals (e.g., “What if surveillance were community-led?”), most models paraphrase prior stances rather than updating them through principled reasoning [84, 85, 86, 87, 88]. These failures reflect a deeper structural problem: we lack benchmarks that probe agents’ causal belief structures, trace their reasoning trajectories, or demand sensitivity to interventions.
@@ 4.2 Theoretical Foundations: Causal, Compositional, Revisable
    Causal: Humans reason in terms of causes and consequences. Even young children exhibit Bayesian-like inference over causal relationships and use interventions to test hypotheses about the world [12, 91]. Mental models are structured around “what caused what”—not just correlation, but explanation. This causal orientation allows for robust generalization and counterfactual reasoning [90].
@@
Together, these three dimensions—causal, compositional, and revisable—form the foundation for what we call reasoning fidelity: the structural integrity of belief formation and revision processes in generative agents.
@@ 4.3 Defining Reasoning Fidelity
Importantly, this architecture is not tied to any particular implementation. While LLMs may serve as one plausible interface for extracting cognitive motifs, the core modeling contribution lies in structuring reasoning as revisable causal graphs—an approach compatible with both symbolic and neural systems [92]. GenMinds exemplifies one such instantiation of this broader modeling principle.
@@
Yet current LLM-based agents fall short of this standard. Most optimize for surface alignment—producing plausible stances like “I support policy X”—without modeling the underlying belief process. They lack persistent belief states, causal coherence, and principled revision under counterfactuals. This results in brittle or contradictory responses, agreement bias between agents, and an absence of traceable justification.
@@ 5.1 GenMinds: A Framework for Modeling Human-Like Reasoning
To build generative agents that simulate human reasoning—not merely output plausible stances—we propose modeling individuals’ internal logic through semi-structured interviews, adaptively conducted by large language models (LLMs). These interviews elicit causal explanations in everyday language (e.g., “why do you support X?” “what does Y influence?”), which are then parsed into directed acyclic graphs representing the participant’s belief structure [3]. Each node encodes a concept (e.g., fairness, safety, family needs), and each edge encodes a directional causal relation, with confidence and polarity scores.
@@
This chain demonstrates the potential of motif-based causal modeling for simulating how real individuals might update beliefs when exposed to policy changes—moving beyond static opinion snapshots.
@@ 5.2 RECAP: Principles for Evaluating Reasoning Fidelity
To advance cognitively aligned simulation, we propose a benchmark framework—RECAP—that shifts evaluation from surface-level correctness to the internal structure and coherence of reasoning.
@@
  • A task—e.g., graph reconstruction, stance explanation, or counterfactual reasoning—that requires structured inference.
@@ 6 A Call for Cognitively Grounded Simulation
As large language models become embedded in social simulations and policy tools, we face a pivotal choice: whether to pursue agents that merely sound human, or agents that can reason in structured, human-like ways. This paper argues for the latter. We call for a shift from behavior-level mimicry to cognitively grounded reasoning—where agents represent beliefs, simulate causal relationships, revise assumptions, and reveal their internal logic.
@@
Adopting reasoning fidelity as a core standard would shift generative agent research from stylistic fluency to structural interpretability. It reshapes alignment evaluation, promotes modular and revisable architectures, and incentivizes cognitively grounded benchmarks. In high-stakes applications—civic simulation, participatory policy, AI governance—agents with causal transparency and revisable beliefs are essential for trust, auditability, and fairness. Without this shift, we risk institutionalizing brittle models that obscure bias and flatten the diversity of public thought.
What Comes Next.

### v3

@@ 3.3 Evaluation Gaps: Structural Limitations of Current Benchmarks
The deficiencies in reasoning fidelity and individuality are compounded by a third layer: evaluation misalignment. Despite the growing sophistication of generative agents, current benchmarks remain optimized for stylistic fluency, local coherence, and plausibility of individual model behavior and output. Most benchmarks treat language as a proxy of thought, implicitly assuming that coherent expression entails coherent reasoning. As a result, current evaluations risk rewarding surface-level coherence while overlooking causal consistency and belief heterogeneity [71]. This first creates what we term as traceability gap: benchmarks assess outputs instead of reasoning trajectories. Stance classification tasks, for example, check whether a model picks a side but remain agnostic about how or why that stance was formed [83, 84]. Dialogue benchmarks reward conversational smoothness, even when agents flip positions over time [85, 86, 87].
The second limitation concerns intervention blindness: most benchmarks assess agents on static inputs but fail to measure their belief revision under counterfactual interventions or other hypothetical perturbations [88, 89, 90, 91, 92]. Finally, current benchmarks leave agents’ positional individuality largely unaddressed. Evaluation suites typically aggregates performance across stances, computing mean accuracy or sentiment agreement but rarely quantify inter-agent divergence or distributional variance. Recent work on pluralistic alignment has begun moving in this direction, measuring whether models maintain epistemic diversity or support multiple internally coherent responses [93, 94, 95]; yet these methods remain limited to small-scale reasoning or dialogue tasks and have not been adopted within social simulation pipelines, where evaluation still focuses on output fluency and stance accuracy.
In sum, evaluation protocols for generative agents remain behaviorally calibrated but structurally ungrounded. Bridging this gap requires a new generation of benchmarks that treat reasoning as a structured process rather than a stylistic performance.
@@ 4.2 Theoretical Foundations: Causal, Compositional, Revisable
    Causal: Humans reason in terms of causes and consequences. Even young children exhibit Bayesian-like inference over causal relationships and use interventions to test hypotheses about the world [16, 98]. Mental models are structured around “what caused what,” emphasizing explanation rather than mere correlation. This causal orientation allows for robust generalization and counterfactual reasoning [97].
@@
Taken together, the three dimensions of causal, compositional, and revisable reasoning form the foundation of what we call reasoning fidelity, defined as the structural integrity of belief formation and revision processes in generative agents.
@@ 4.3 Defining Reasoning Fidelity
Importantly, this architecture is not tied to any particular implementation. While LLMs may serve as one plausible interface for extracting cognitive motifs, the core modeling contribution lies in structuring reasoning as revisable causal graphs. This approach is compatible with both symbolic and neural systems [99], and GenMinds exemplifies one such instantiation of this broader modeling principle.
@@
Yet current LLM-based agents fall short of this standard. Most optimize for surface alignment by producing plausible stances such as “I support policy X,” without modeling the underlying belief process. They lack persistent belief states, causal coherence, and principled revision under counterfactuals. This results in brittle or contradictory responses, agreement bias between agents, and an absence of traceable justification.
@@ 5.1 GenMinds: A Framework for Modeling Human-Like Reasoning
To build generative agents that simulate human reasoning rather than merely output plausible stances, we propose modeling individuals’ internal logic through semi-structured interviews, adaptively conducted by large language models (LLMs). These interviews elicit causal explanations in everyday language (e.g., “why do you support X?” “what does Y influence?”), which are then parsed into directed acyclic graphs representing the participant’s belief structure [3]. Each node encodes a concept (e.g., fairness, safety, family needs), and each edge encodes a directional causal relation, with confidence and polarity scores.
@@
This chain demonstrates the potential of motif-based causal modeling to simulate how real individuals update their beliefs in response to policy changes, thereby moving beyond static opinion snapshots.
@@ 5.2 RECAP: Principles for Evaluating Reasoning Fidelity
To advance cognitively aligned simulation, we propose a benchmark framework called RECAP, that shifts evaluation from surface-level correctness to the internal structure and coherence of reasoning.
@@
  • A task such as graph reconstruction, stance explanation, or counterfactual reasoning that requires structured inference.
@@ 6 A Call for Cognitively Grounded Simulation
As large language models become embedded in social simulations and policy tools, we face a pivotal choice: whether to pursue agents that merely sound human, or agents that can reason in structured, human-like ways. This paper argues for the latter. We call for a shift from behavior-level mimicry to cognitively grounded reasoning, where agents represent beliefs, simulate causal relationships, revise assumptions, and reveal their internal logic.
@@
Adopting reasoning fidelity as a core standard would shift generative agent research from stylistic fluency to structural interpretability. It reshapes alignment evaluation, promotes modular and revisable architectures, and incentivizes cognitively grounded benchmarks. In high-stakes applications such as civic simulation, participatory policy, and AI governance, agents with causal transparency and revisable beliefs are essential for trust, auditability, and fairness. Without this shift, we risk institutionalizing brittle models that obscure bias and flatten the diversity of public thought.
Open Challenges and Next Steps
@@
Alongside these efforts, we identify several open challenges:
  • Constructing causal belief networks from natural language transcripts remains challenging, due to ambiguity in concept identification, causal direction, polarity, and conceptual granularity;
  • Causality alone cannot capture the full range of human reasoning. People also rely on associative, analogical, and emotional processes that resist strict symbolic modeling. Our initial focus on casuality is a strategic and computationally tractable starting point, not an endpoint.
@@
Acknowledgments
This work builds on research originally developed at MIT. We thank Kent Larson, Jinhua Zhao, Paul Liang, and Deb Roy for their valuable guidance and discussions throughout this work. We are especially grateful to Kent Larson for his continuous mentorship, to Deb Roy for his conceptual insights and high standards that shaped the framing of this work, to Jinhua Zhao for his thoughtful resonance and encouragement, and to Paul Liang for providing key technical perspectives.